OpenStudy (anonymous):

Find the length of the side of the square 9x^2 + 6x +1

OpenStudy (anonymous):

\[L=\sqrt{1-(df/dx})^2 so friest take the dervetive of the equation you have then squre both side u get the length\]

OpenStudy (anonymous):

f`(x)=18X+6 F`(X)^2=(18X+6)^2 L=\[\sqrt{1-(18x+6)^2}\]

OpenStudy (anonymous):

i hope u understand it

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

is it clear

OpenStudy (anonymous):

Ya thanks